A sample of nitrogen, N2, occupies 45.0 mL at 27 °C and 600 torr. What pressure will it have if cooled to –73 °C while the volume remains constant?

1
Expert's answer
2021-12-20T17:08:04-0500

"V = constant = 45.0 \;mL \\

T_1 = 27+273 = 300 \;K \\

p_1 = 600 \; torr \\

T_2 = -73 + 273 = 200 \;K \\

\frac{p_1}{T_1} = \frac{p_2}{T_2} \\

p_2 = \frac{p_1 T_2}{T_1} \\

= \frac{600 \times 200}{300} = 400 \; torr"

Answer: 400 torr